# Unifying Multiple Foundation Models for Advanced Computational Pathology

### v3

@@ 4 Patch-level Classification
where $\lambda_{\mathrm{distill}}=0.01$. The distillation term is composed of a cosine-similarity loss and a Huber loss applied at each semantic scale. To properly handle class imbalance, we report Balanced Accuracy, Weighted F1, and AUC as the evaluation metrics.
@@
This dataset comprises 101,398 H&E-stained image patches, also referred to as regions of interest (ROIs), each measuring 512 × 512 pixels at a resolution of 0.48 microns per pixel (mpp). These ROIs were extracted from 200 formalin-fixed paraffin-embedded (FFPE) whole-slide images (WSIs) of colorectal biopsies[pataki2022huncrc]. Designed to support the training of machine learning models for colorectal tissue classification, the dataset includes annotations for nine classes: Adenocarcinoma (4,315 ROIs), High-Grade Dysplasia (2,281 ROIs), Low-Grade Dysplasia (55,787 ROIs), Inflammation (763 ROIs), Tumor Necrosis (365 ROIs), Suspicious for Invasion (570 ROIs), Resection Edge (534 ROIs), Technical Artifacts (3,470 ROIs), and Normal Tissue (31,323 ROIs). For training and evaluation, the dataset follows the official split, with 76,753 ROIs used for training, 11,327 ROIs for validation, and 11,328 ROIs for testing. The experimental results for this dataset are reported in Supplementary Table 4.
@@
This dataset comprises 9,536 H&E-stained image patches (ROIs) extracted from 292 whole-slide images (WSIs)[barbano2021unitopatho]. Its primary purpose is to facilitate the training of deep neural networks for colorectal polyp classification and adenoma grading. The dataset includes annotations for six classes: Normal Tissue (950 ROIs), Hyperplastic Polyp (545 ROIs), Tubular Adenoma with High-Grade Dysplasia (454 ROIs), Tubular Adenoma with Low-Grade Dysplasia (3,618 ROIs), Tubulo-Villous Adenoma with High-Grade Dysplasia (916 ROIs), and Tubulo-Villous Adenoma with Low-Grade Dysplasia (2,186 ROIs). For training and evaluation, the dataset follows the official split, with 6,270 ROIs used for training, 1,199 ROIs for validation, and 1,200 ROIs for testing. The experimental results for this dataset are reported in Supplementary Table 5.
@@
PCAM consists of 327,680 color images with a resolution of $96\times 96$ pixels, extracted from the CAMELYON16 dataset[bejnordi2017diagnostic][veeling2018rotation]. Each image is annotated with a binary label indicating the presence of metastatic tissue. For training and evaluation, we adopt the official train-validation-test split of 262,144, 32,768, and 32,768 ROIs, respectively, and resize all images to $224\times 224$ pixels. The experimental results for this dataset are reported in Supplementary Table 6.
@@
ESCA contains 367,229 patches, each with a resolution of $256\times 256$ pixels, sourced from various institutions: University Hospital Cologne (UKK), Landesklinikum Wiener Neustadt (WNS), TCGA, and University Hospital Berlin Charité (CHA)[tolkach2023artificial]. These patches were annotated into 11 classes, including adventitia, lamina propria mucosae, muscularis mucosae, muscularis propria, regression tissue, mucosa gastric, mucosa esophagus, submucosa, submucosal glands, tumor, and ulceration. For training and evaluation, we use the CHA dataset (178,187 ROIs) as the training set and combine UKK, WNS, and TCGA (189,142 ROIs) for testing. All images were resized to $224\times 224$ pixels. The experimental results for this dataset are reported in Supplementary Table 7.
@@
PanCancer-TIL consists of 304,097 images with a resolution of $100\times 100$ pixels, labeled as TIL-positive (54,910 ROIs) and TIL-negative (249,187 ROIs)[saltz2018spatial]. The official train-validation-test split of 209,221, 38,601, and 56,275 ROIs is adopted, and all images are resized to $256\times 256$ pixels. The experimental results for this dataset are reported in Supplementary Table 8.
@@
The CRC-100K dataset is composed of the NCT-CRC-HE-100K and CRC-VAL-HE-7K cohorts [kather2019predicting]. NCT-CRC-HE-100K contains 100,000 non-overlapping $224\times 224$ H&E patches sourced from 86 colorectal cancer whole-slide images obtained from the NCT biobank and the UMM pathology archive. The CRC-VAL-HE-7K subset includes 7,180 $224\times 224$ patches collected from 50 patients diagnosed with colorectal adenocarcinoma. The dataset covers nine histological tissue categories: adipose (ADI, 11,745 ROIs), background (BACK, 11,413), debris (DEB, 11,851), lymphocytes (LYM, 12,191), mucus (MUC, 9,931), smooth muscle (MUS, 14,128), normal colon mucosa (NORM, 9,504), cancer-associated stroma (STR, 10,867), and colorectal adenocarcinoma epithelium (TUM, 15,550). We follow the official train–test split of 100,000 and 7,180 patches, respectively. The experimental results for this dataset are reported in Supplementary Table 9.
@@
This dataset consists of 52,713 regions of interest (ROIs), each with a resolution of $300\times 300$ pixels, from TCGA-KIRC and Helsinki [brummer2023computational]. The dataset includes six categories: renal cancer (cancer, 13,057 ROIs), normal renal tissue (normal, 8,652 ROIs), stromal tissue (stroma, 5,460 ROIs), red blood cells (blood, 996 ROIs), empty background (empty, 16,026 ROIs), and other textures (other, 8,522 ROIs). For training and evaluation, we focus on four specific categories: cancer, stroma, normal, and blood, excluding the ”empty” and ”other” categories. The samples are shuffled and split using a 22,530:5,635 ratio for train/test. The experimental results for this dataset are reported in Supplementary Table 10.
@@
The BACH dataset consists of 400 images ($2048\times 1536$ pixels) labeled into four classes: Normal (100 ROIs), Benign (100 ROIs), In situ carcinoma (100 ROIs), and Invasive carcinoma (100 ROIs)[aresta2019bach]. For training and evaluation, all ROIs are resized to $224\times 224$ pixels, and a label-stratified train-test split of 320:80 ROIs is used. The experimental results for this dataset are reported in Supplementary Table 11.
@@
The CRC-MSI dataset includes 51,918 histology images with a resolution of $512\times 512$ pixels, obtained from the TCGA database [kather2019deep]. The images are annotated with MSI status, where patients are classified as MSI-H (high MSI) or NonMSIH (either MSI-L or MSS). For training and evaluation, we use the official 19,557:32,361 train-test split. The experimental results for this dataset are reported in Supplementary Table 12.
@@
The PanCancer-TCGA dataset contains 271,170 images with dimensions $256\times 256$ pixels, extracted from 8,736 TCGA WSIs across 32 different cancer types [komura2022universal]. The images are annotated with 32 classes, including Head and Neck Squamous Cell Carcinoma (11,790 ROIs), Bladder Urothelial Carcinoma (9,990 ROIs), Lung Squamous Cell Carcinoma (16,560 ROIs), Glioblastoma Multiforme (23,740 ROIs), among others. For training and evaluation, we use a 21,736:54,342 train/test split. The experimental results for this dataset are reported in Supplementary Table 13.
@@
The Chaoyang dataset includes colon tissue patches from Chaoyang hospital, comprising 1,816 normal ROIs, 1,163 serrated ROIs, 2,244 adenocarcinoma ROIs, and 937 adenoma ROIs [zhu2021hard]. For training and evaluation, all patches are resized to $224\times 224$ pixels, and the official train-test split (4,021:2,139 ROIs) is used. The experimental results for this dataset are reported in Supplementary Table 14.
5 Survival Analysis
Survival analysis has long been a standard approach for modeling time-to-event data in oncology, focusing on outcomes such as disease progression or overall survival. When applied to whole-slide images (WSIs), it provides new opportunities to explore tissue-level patterns and predict patient outcomes. By integrating survival analysis with WSIs, researchers can examine correlations between specific morphological characteristics and clinical prognosis.
[…]
In this study, we adopt the Attention-Based Multiple Instance Learning (ABMIL) framework [ilse2018attention] for survival analysis, employing the Negative Log-Likelihood (NLL) loss. we introduce a distillation loss term for model compression. Specifically, we employ Shazam student model, distilled from a larger teacher model, using a weighted distillation loss with $\lambda=0.01$ to align the student’s predictions with the teacher’s. Optimization is performed using AdamW with a learning rate of $2\times 10^{-4}$ and weight decay of $1\times 10^{-3}$.The dataset partitioning strategy follows that described in the GPFM study [ma2024towards]. To assess the impact of different pathology foundation models on survival prediction, we perform 5-fold cross-validation and use the C-index as the evaluation metric.
Experiments are conducted on 10 TCGA datasets: TCGA-BRCA, TCGA-BLCA, TCGA-KIRC, TCGA-LUAD, TCGA-STAD, TCGA-LUSC, TCGA-COADREAD, TCGA-GBMLGG, TCGA-SKCM, and TCGA-CESC. The results are reported in Supplementary Data Table 4- 7.
Breast Cancer Survival Prediction.
We leverage the TCGA-BRCA dataset[tcga-dataset] for breast cancer survival prediction, which consists of 1,023 cases (1,089 WSIs). Following the GPFM partitioning strategy, we maintain uniform censorship between the training and testing sets and apply stratified binning of survival times into four equally sized bins. For each fold of the 5-fold cross-validation, we perform an 8:2 label-stratified split, resulting in 821 training cases and 202 testing cases. The mean C-index and standard deviation across the folds are presented in Extended Data Table A16.
Bladder Cancer Survival Prediction.
For bladder cancer survival prediction, we use the TCGA-BLCA dataset[tcga-dataset], containing 376 cases (446 WSIs). Using the same stratified binning and partitioning strategy as BRCA, each fold allocates 305 cases to training and 71 cases to testing. The results are shown in the BLCA column of Extended Data Table A16.
Kidney Renal Clear Cell Carcinoma Survival Prediction.
The TCGA-KIRC dataset[tcga-dataset] comprises 498 cases (504 WSIs). Following the same partitioning strategy, each fold assigns 401 cases to training and 97 cases to testing. Results are reported in the KIRC column of Extended Data Table A16.
Stomach Adenocarcinoma Survival Prediction.
The TCGA-STAD dataset[tcga-dataset] contains 363 cases (389 WSIs). In each fold, 293 cases are allocated to training and 70 to testing using the same stratified splitting procedure. Results are shown in the STAD column of Extended Data Table A17.
The TCGA-CESC[tcga-dataset] dataset includes 250 cases (260 WSIs). Using the same strategy, we assign 203 cases to training and 47 cases to testing per fold. Results are presented in the CESC column of Extended Data Table A17.
Lung Adenocarcinoma Survival Prediction.
For TCGA-LUAD[tcga-dataset], we use 455 cases (518 WSIs), splitting each fold into 366 training cases and 89 testing cases. Results for both datasets are shown in Extended Data Table A18.
Lung Squamous Cell Carcinoma Survival Prediction.
The TCGA-LUSC dataset[tcga-dataset] comprises 452 cases (484 WSIs). Using the same stratified splitting approach, 365 cases are assigned to training and 87 to testing in each fold. Results are reported in Extended Data Table A18.
Colon & Rectum Adenocarcinoma Survival Prediction.
We combine TCGA-COAD[tcga-dataset] (426 cases, 434 WSIs) and TCGA-READ (153 cases, 154 WSIs) into a single dataset of 579 cases (588 WSIs). Each fold assigns 464 cases to training and 115 to testing. Results are shown in Extended Data Table A19.
Glioma Survival Prediction.
We merge TCGA-GBM (370 cases, 823 WSIs) and TCGA-LGG[tcga-dataset] (460 cases, 778 WSIs) into a combined glioma dataset of 830 cases (1,601 WSIs). Per fold, 667 cases are used for training and 163 for testing. Results are in Extended Data Table A19.
Skin Cutaneous Melanoma Survival Prediction.
The TCGA-SKCM dataset[tcga-dataset] contains 415 cases (456 WSIs). Each fold uses 337 cases for training and 78 for testing. Results are shown in Extended Data Table A19.
@@ 6 Pathology VQA
We evaluate our pathology foundation models on the PathVQA benchmark [he2020pathvqa], a standard dataset for assessing visual question answering (VQA) capabilities in pathology. PathVQA requires models to jointly interpret visual content and natural-language questions, providing a comprehensive assessment of multimodal medical reasoning. We adopt the Multi-modal Unified Medical Captioning (MUMC) framework [mumc], which demonstrates strong performance on PathVQA by tightly integrating textual and visual representations.
@@
The experimental results are summarized in Supplementary Table 18. Integrating Shazam as a multi-level visual expert leads to a substantial improvement in accuracy, demonstrating the effectiveness of hierarchical feature integration for supporting fine-grained multimodal reasoning in pathology.
@@
All experiments were conducted in the Python environment (Python 3.10.0). Model training and inference were performed using PyTorch (version 2.8.0, compiled with CUDA 12.8), together with TorchVision 0.23.0. Downstream machine learning experiments used Scikit-learn 1.7.2 and NumPy 2.2.6. Whole-slide image (WSI) processing relied on the OpenSlide-Python interface (version 1.4.2) as well as the CLAM [lu2021data] codebase. For WSI processing, we relied on openslide-python (version 1.2.0) and the CLAM [lu2021data] codebase. Pathology VQA evaluation was conducted using the MUMC [mumc] codebase. All computations were executed on a workstation equipped with 8 $\times$ NVIDIA GeForce RTX 4090 GPUs.
@@
The HEST-1k dataset used in this study is publicly available at Hugging Face (https://huggingface.co/datasets/MahmoodLab/hest). Processed data and model outputs generated during this study are available from the corresponding author upon reasonable request.
The Survival Analysis tasks are based on clinical and genomic data that can be accessed through the GDC Data Portal (https://portal.gdc.cancer.gov/).
The PathVQA dataset used for visual question answering experiments is available at Github (https://github.com/UCSD-AI4H/PathVQA).
All tile-level classification tasks in this study are conducted on publicly available histopathology datasets, including:
@@
The code of the Shazam have been made available on GitHub https://github.com/Tuner12/Shazam

### v4

@@ 5 Tile-level Classification
where $\lambda_{\mathrm{distill}}=0.01$. The distillation term is composed of a cosine-similarity loss and a Huber loss applied at each semantic scale. To properly handle class imbalance, we report Balanced Accuracy, Weighted F1, and Top1-ACC as the evaluation metrics.
@@
This dataset comprises 101,398 H&E-stained image patches, also referred to as regions of interest (ROIs), each measuring 512 × 512 pixels at a resolution of 0.48 microns per pixel (mpp). These ROIs were extracted from 200 formalin-fixed paraffin-embedded (FFPE) whole-slide images (WSIs) of colorectal biopsies[35]. Designed to support the training of machine learning models for colorectal tissue classification, the dataset includes annotations for nine classes: Adenocarcinoma (4,315 ROIs), High-Grade Dysplasia (2,281 ROIs), Low-Grade Dysplasia (55,787 ROIs), Inflammation (763 ROIs), Tumor Necrosis (365 ROIs), Suspicious for Invasion (570 ROIs), Resection Edge (534 ROIs), Technical Artifacts (3,470 ROIs), and Normal Tissue (31,323 ROIs). For training and evaluation, the dataset follows the official split, with 76,753 ROIs used for training, 11,327 ROIs for validation, and 11,328 ROIs for testing. The experimental results for this dataset are reported in Supplementary Table 8.
@@
This dataset comprises 9,536 H&E-stained image patches (ROIs) extracted from 292 whole-slide images (WSIs)[6]. Its primary purpose is to facilitate the training of deep neural networks for colorectal polyp classification and adenoma grading. The dataset includes annotations for six classes: Normal Tissue (950 ROIs), Hyperplastic Polyp (545 ROIs), Tubular Adenoma with High-Grade Dysplasia (454 ROIs), Tubular Adenoma with Low-Grade Dysplasia (3,618 ROIs), Tubulo-Villous Adenoma with High-Grade Dysplasia (916 ROIs), and Tubulo-Villous Adenoma with Low-Grade Dysplasia (2,186 ROIs). For training and evaluation, the dataset follows the official split, with 6,270 ROIs used for training, 1,199 ROIs for validation, and 1,200 ROIs for testing. The experimental results for this dataset are reported in Supplementary Table 9.
@@
Derived from the CAMELYON16 [8][44], the PCAM dataset comprises 327,680 color histopathology patches, each with a native resolution of $96\times 96$ pixels. The dataset poses a binary classification problem, where each image is labeled according to the presence of metastatic tissue. In our experiments, we adhere to the standard data partition, utilizing 262,144 images for training and 32,768 images each for validation and testing. Prior to model input, all regions of interest (ROIs) are upscaled to $224\times 224$ pixels. The experimental results for this dataset are reported in Supplementary Table 10.
@@
ESCA contains 367,229 patches, each with a resolution of $256\times 256$ pixels, sourced from various institutions: University Hospital Cologne (UKK), Landesklinikum Wiener Neustadt (WNS), TCGA, and University Hospital Berlin Charité (CHA)[41]. These patches were annotated into 11 classes, including adventitia, lamina propria mucosae, muscularis mucosae, muscularis propria, regression tissue, mucosa gastric, mucosa esophagus, submucosa, submucosal glands, tumor, and ulceration. For training and evaluation, we use the CHA dataset (178,187 ROIs) as the training set and combine UKK, WNS, and TCGA (189,142 ROIs) for testing. All images were resized to $224\times 224$ pixels. The experimental results for this dataset are reported in Supplementary Table 11.
@@
PanCancer-TIL consists of 304,097 images with a resolution of $100\times 100$ pixels, labeled as TIL-positive (54,910 ROIs) and TIL-negative (249,187 ROIs)[37]. The official train-validation-test split of 209,221, 38,601, and 56,275 ROIs is adopted, and all images are resized to $256\times 256$ pixels. The experimental results for this dataset are reported in Supplementary Table 12.
@@
The CRC-100K dataset is composed of the NCT-CRC-HE-100K and CRC-VAL-HE-7K cohorts [23]. NCT-CRC-HE-100K contains 100,000 non-overlapping $224\times 224$ H&E patches sourced from 86 colorectal cancer whole-slide images obtained from the NCT biobank and the UMM pathology archive. The CRC-VAL-HE-7K subset includes 7,180 $224\times 224$ patches collected from 50 patients diagnosed with colorectal adenocarcinoma. The dataset covers nine histological tissue categories: adipose (ADI, 11,745 ROIs), background (BACK, 11,413), debris (DEB, 11,851), lymphocytes (LYM, 12,191), mucus (MUC, 9,931), smooth muscle (MUS, 14,128), normal colon mucosa (NORM, 9,504), cancer-associated stroma (STR, 10,867), and colorectal adenocarcinoma epithelium (TUM, 15,550). We follow the official train and test split of 100,000 and 7,180 patches, respectively. The experimental results for this dataset are reported in Supplementary Table 13.
@@
We utilize a dataset of 52,713 regions of interest (ROIs) sourced from the TCGA-KIRC and Helsinki cohorts [10], with each patch standardized to $300\times 300$ pixels. While the original dataset spans six categories, we restrict our analysis to four clinically relevant classes: renal cancer (13,057), normal renal tissue (8,652), stroma (5,460), and red blood cells (996). After excluding the ’empty’ (16,026) and ’other’ (8,522) categories, the remaining 28,165 samples are shuffled and partitioned into a training set of 22,530 and a test set of 5,635. The experimental results for this dataset are reported in Supplementary Table 14.
@@
The BACH dataset [3] comprises 400 high-resolution histology images ($2048\times 1536$), evenly distributed across four classes: Normal, Benign, In situ carcinoma, and Invasive carcinoma (100 samples per class). To facilitate model training, we resize all images to $224\times 224$ pixels and employ a stratified random split, allocating 320 images for training and 80 for evaluation. The experimental results for this dataset are reported in Supplementary Table 15.
@@
The CRC-MSI dataset[24] consists of 51,918 histological patches ($512\times 512$ pixels) derived from the TCGA colorectal cancer cohort. The data is categorized by Microsatellite Instability (MSI) status, distinguishing between patients with high instability (MSI-H) and those without (Non-MSI-H, comprising MSI-L and MSS). We adhere to the official data partition, utilizing 19,557 images for training and 32,361 for testing. The experimental results for this dataset are reported in Supplementary Table 16.
@@
The PanCancer-TCGA dataset [25] is a large-scale collection of 271,170 histological patches ($256\times 256$ pixels) derived from 8,736 Whole Slide Images (WSIs) in the TCGA archive. Spanning 32 distinct cancer subtypes, the dataset includes major classes such as Head and Neck Squamous Cell Carcinoma (11,790 ROIs), Bladder Urothelial Carcinoma (9,990 ROIs), Lung Squamous Cell Carcinoma (16,560 ROIs), and Glioblastoma Multiforme (23,740 ROIs). For training and evaluation, we use a 21,736:54,342 train/test split. The experimental results for this dataset are reported in Supplementary Table 17.
@@
Sourced from Chaoyang Hospital [49], the Chaoyang dataset consists of 6,160 colon tissue tiles categorized into four distinct classes: Normal (1,816), Serrated (1,163), Adenocarcinoma (2,244), and Adenoma (937). We adhere to the standard evaluation protocol, resizing all images to $224\times 224$ pixels and employing the official training and testing partition of 4,021 and 2,139 samples, respectively. The experimental results for this dataset are reported in Supplementary Table 18.
@@ 6 Pathology VQA
We evaluate our pathology foundation models on the PathVQA benchmark [19], a standard dataset for assessing visual question answering (VQA) capabilities in pathology. PathVQA requires models to jointly interpret visual content and natural-language questions, providing a comprehensive assessment of multimodal medical reasoning. We adopt the Multi-modal Unified Medical Captioning (MUMC) framework [26], which demonstrates strong performance on PathVQA by tightly integrating textual and visual representations.
@@
The experimental results are summarized in Supplementary Table 19. Integrating Shazam as a multi-level visual expert leads to a substantial improvement in accuracy, demonstrating the effectiveness of hierarchical feature integration for supporting fine-grained multimodal reasoning in pathology
7 Ablation Study
7.1 Spatial Transcriptomic Ablation Study
We conducted an ablation study under 5-fold cross-validation by sequentially removing teacher models from the ensemble according to their standalone performance ranking. We first established an ordered list of teacher feature extractors based on single-teacher validation results, then initialized the ablation sequence from the complete teacher set used in Shazam and generated a series of reduced-teacher configurations by removing the top-ranked teacher, followed by the next-ranked teacher, iteratively until only a minimal subset remained. Throughout all ablation runs, we held the data splits, model architecture, training pipeline, and evaluation protocol fixed to ensure that observed differences across configurations can be attributed solely to the removal of specific teachers. The results are reported in Supplementary Table 20.
7.2 Survival Analysis Ablation Study
For WSI-level survival prediction, we conducted an ablation study on the TCGA-COADREAD cohort to assess the contribution of hierarchical multi-level features. The full Shazam model integrates low-, middle-, and high-level representations extracted from each teacher model. To isolate the effect of feature depth, we evaluated reduced variants that use only low-level, mid-level, and high-level features or paired combinations of {low, mid}, {low, high}, or {mid, high}. To ensure a controlled comparison, all variants were trained under the same data splits, optimization configuration, and student-model design. Thus, changes in performance directly reflect the contribution of the feature-level combinations being tested. A complete comparison of all feature-level variants is summarized in Supplementary Table 21.
7.3 Tile-level Ablation Study
To quantify the contribution of the Mixture of Experts (MoE) module in tile-level classification, we conducted an ablation in which the MoE fusion was removed and the five teacher features were directly stacked without adaptive gating. All other aspects of the training pipeline—including data partitions, optimization scheme, and student model architecture—were kept unchanged, ensuring that any performance difference arises specifically from the removal of the MoE fusion mechanism. By comparing the full MoE-enabled model with its MoE-ablated counterpart on the PCAM and CCRCC datasets, we assess how much adaptive multi-teacher fusion improves tile-level prediction performance. Detailed results for this ablation are provided in Supplementary Table 22 and 23.
@@ 7.3 Tile-level Ablation Study
All experiments were conducted in the Python environment (Python 3.10.0). Model training and inference were performed using PyTorch (version 2.8.0, compiled with CUDA 12.8), together with TorchVision 0.23.0. Downstream machine learning experiments used Scikit-learn 1.7.2 and NumPy 2.2.6. Whole-slide image (WSI) processing relied on the OpenSlide-Python interface (version 1.4.2) as well as the CLAM [27] codebase. Pathology VQA evaluation was conducted using the MUMC [26] codebase. All computations were executed on a workstation equipped with 8 $\times$ NVIDIA GeForce RTX 4090 GPUs.
@@
The HEST-1K dataset used in this study is publicly available through Hugging Face at https://huggingface.co/datasets/MahmoodLab/hest. Clinical and genomic data for survival analysis experiments can be accessed via the Genomic Data Commons (GDC) Data Portal (https://portal.gdc.cancer.gov/). The PathVQA dataset used for visual question answering experiments is available at https://github.com/UCSD-AI4H/PathVQA.
Processed data and model outputs generated during this study are available from the corresponding author upon reasonable request.
@@
The code of the Shazam is available on GitHub https://github.com/Tuner12/Shazam